OpenStudy (anonymous):

(a^5b^-7)(a^-4b^9) trying to check my answer

OpenStudy (anonymous):

do you multiply the exponents? and leave it at a^-20b^-63

jimthompson5910 (jim_thompson5910):

No you add the corresponding exponents

OpenStudy (pfenn1):

\[a^xa^y=a^{x+y}\]

jimthompson5910 (jim_thompson5910):

ex: a^5 times a^-4 = a^(5 + (-4) ) = a^(5 - 4) = a^1 = a So a^5 times a^-4 = a

OpenStudy (anonymous):

so thatll be ab^2

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thanks, are you good with dividing these?

OpenStudy (pfenn1):

Yes

jimthompson5910 (jim_thompson5910):

dividing is the opposite: you subtract corresponding exponents

OpenStudy (anonymous):

(2a^-3b^2)^-1 --------- (6a^2b^4)

jimthompson5910 (jim_thompson5910):

The first thing to do is to simplify (2a^-3b^2)^-1 Do you know how to do that?

OpenStudy (anonymous):

add -1 to the exponents?

OpenStudy (anonymous):

2a^-4b?

jimthompson5910 (jim_thompson5910):

No you first write 2 as 2^1 Then you multiply the outer exponent of -1 by every exponent So (2a^-3b^2)^-1 (2^1a^-3b^2)^-1 2^(1*-1) * a^(-3*-1) * b^(2*-1) 2^(-1) a^3 b^(-2) This means that (2a^-3b^2)^-1 simplifies to 2^(-1) a^3 b^(-2)

OpenStudy (anonymous):

it looks really confusing on the computer screen with all the ^'s

jimthompson5910 (jim_thompson5910):

wait a second, was that -1 outside the entire group or just outside the numerator?

OpenStudy (anonymous):

entire group

jimthompson5910 (jim_thompson5910):

is the original problem \[\Large \left(\frac{2a^{-3}b^{2}}{6a^2b^4}\right)^{-1}\] OR... is it \[\Large \frac{\left(2a^{-3}b^{2}\right)^{-1}}{6a^2b^4}\]

OpenStudy (anonymous):

first one

jimthompson5910 (jim_thompson5910):

alright thx for clarifying, one sec

OpenStudy (anonymous):

sorry i shouldve wrote it a little better

jimthompson5910 (jim_thompson5910):

no worries

jimthompson5910 (jim_thompson5910):

Here's how you tackle this problem... \[\Large \left(\frac{2a^{-3}b^{2}}{6a^2b^4}\right)^{-1}\] \[\Large \left(\left(\frac{2}{6}\right)\left(\frac{a^{-3}}{a^2}\right)\left(\frac{b^{2}}{b^4}\right)\right)^{-1}\] \[\Large \left(\left(\frac{1}{3}\right)\left(\frac{a^{-3}}{a^2}\right)\left(\frac{b^{2}}{b^4}\right)\right)^{-1}\] \[\Large \left(\left(\frac{1}{3}\right)a^{-3-2}*b^{2-4}\right)^{-1}\] \[\Large \left(\left(\frac{1}{3}\right)a^{-5}*b^{-2}\right)^{-1}\] \[\Large \left(\left(\frac{1}{3}\right)\left(\frac{1}{a^5}\right)\left(\frac{1}{b^2}\right)\right)^{-1}\] \[\Large \left(\frac{1}{3a^5b^2}\right)^{-1}\] \[\Large \left(\frac{3a^5b^2}{1}\right)^{1}\] \[\Large 3a^5b^2\] So \[\Large \left(\frac{2a^{-3}b^{2}}{6a^2b^4}\right)^{-1}\] fully simplifies to \[\Large 3a^5b^2\]

jimthompson5910 (jim_thompson5910):

The idea in the second to last step is that when you have a fraction raised to a negative exponent, you flip the fraction to make the exponent positive.

OpenStudy (anonymous):

thanks bro helped alot!!

jimthompson5910 (jim_thompson5910):

you're welcome